pl.

( \begin{matrix}t & z \\ \overline{z} & s \end{matrix} )

2 × 2-es önadjungált mátrix, ahol t, s ∈ ℝ, z ∈ C.

\underbrace{x_{0} · ( \begin{matrix}1 & 0 \\ 0 & 1 \end{matrix} ) + x_{1} · ( \begin{matrix}0 & 1 \\ 1 & 0 \end{matrix} ) + x_{2} · ( \begin{matrix}0 & −i \\ i & 0 \end{matrix} ) + x_{3} · ( \begin{matrix}1 & 0 \\ 0 & −1 \end{matrix} )}_{Pauli-mátrixok} = ( \begin{matrix}x_{0} + x_{3} & x_{1} − i · x_{2} \\ x_{1} + i · x_{2} & x_{0} − x_{3} \end{matrix} )

V_{1}, V_{2} vektorterek

1. Definíció. φ: V_{1} → V_{2} leképezés. φ lineáris, ha φ(λf + μg) = λ · φ(f) + μ · φ(g), ahol f, g ∈ V_{1}, λ, μ ∈ ℂ. Ha V_{2} = ℂ, akkor φ lineáris funkcionál.